Is there some way someone can help me with this? I get lost after the first sentence.
A disk jockey must play 14 commercial spots during one hour of a radio show. Each commercial is either 30 seconds or 60 seconds long. If the total commercial time during 1 hour is 11 minutes, how many 60 second commercials were played that hour. I thank you.very much, kate

Answer by RAY100(1637)

let 14 = total comm spots
let x = # of 30 sec spots, therefore 14-x = # of 60 sec spots
total comm time is 11 min = 11 *60 sec = 660 sec
total time = 30 x + 60 ( 14-x) = 660
simplifying 30 x -60 x + 840 = 660
-30x = - 180
x=6 ( # of 30 sec spots)
and (14-x) = 14-6 = 8 (# 0f 60 sec spots)
check (6) (30) + (8) (60) = 660 sec = 11 min ok